OpenStudy (anonymous):

how to find the recursive rule for 3,4,1,-3,-4...?

OpenStudy (freckles):

do you see any kind of pattern?

OpenStudy (anonymous):

It's neither arithmetic or geometric

OpenStudy (freckles):

like how can you get 1 from the first two numbers? how can you get -3 from the previous two numbers?

OpenStudy (anonymous):

I couldn't find a common difference or a common rate, so i couldn't find the rule

OpenStudy (freckles):

there is no common difference or ratio

OpenStudy (anonymous):

could somebody help me?

OpenStudy (freckles):

do you know that 4-3=1? and that 1-4=-3?

OpenStudy (anonymous):

Yes, I did that. so it's +1,-3,-4, -1

OpenStudy (freckles):

3,4,1,-3,-4 notice 4-3=1 1-4=-3 -3-1=-4 and so on... notice you are taking the second term in every pair and subtracting the first number from that pair

OpenStudy (freckles):

so you have term-previous term=next term

OpenStudy (freckles):

try setting up the recurrence relation from there

OpenStudy (anonymous):

so I find the rule for 1, -3, and -4 first?

OpenStudy (freckles):

i don't know what you mean i already gave you the recurrence relation you just have to put the appropriate math symbols in

OpenStudy (freckles):

\[\text{ notice you have } \\ a_2-a_1=a_3 \\ a_3-a_2=a_4 \\ \\ \cdots \]

OpenStudy (freckles):

so ... \[a_n-?=?\] what are the question marks in this equation

OpenStudy (anonymous):

ohh, so its the previous term subtracted from the previous previous term

OpenStudy (freckles):

it is what I said earlier term-previous term=next term

OpenStudy (anonymous):

a(n)=a(n-2)-a(n-1)

OpenStudy (anonymous):

Thanks! I thought it was going to be a quadratic recursive rule, so i was confused.

OpenStudy (freckles):

your equation says term=previous previous term-previous term so testing... plug in n=3 \[a_3=a_{3-2}-a_{3-1} \\ a_3=a_1-a_2 \\ 1=3-4 \\ 1=-1 \\ \text{ but this is \not true }\] maybe you meant term=previous term-previous previous term

OpenStudy (freckles):

if you have a term a(n) then the previous term is a(n-1) and the previous previous term is a(n-2)

OpenStudy (freckles):

and if you wanted to use my equation you would need to find the next term to a(n) which would be a(n+1)

OpenStudy (anonymous):

a(n)=a(n-1)-a(n-2)?

OpenStudy (freckles):

yes that looks better

OpenStudy (anonymous):

Thank you again for your help!

OpenStudy (freckles):

np I think I see what you were trying to do. You were trying to look at first maybe even second differences.

OpenStudy (freckles):

The first thing I thought of and since it say recurrence relation (and since it had no common ratio or common difference) I try to look for a pattern in terms of previous or next or both types of terms since that is how you write a recurrence relation. could you write a recurrence relation for the following? 1,1,2,3,5,8,...

OpenStudy (anonymous):

a(n)=a(n-2)+a(n-1)

OpenStudy (freckles):

oops sorry my notifications didn't work

OpenStudy (freckles):

that is right lol did you know that was the Fibonacci sequence when you answered?

OpenStudy (anonymous):

lol yep.

OpenStudy (freckles):

such a famous sequence :p

OpenStudy (anonymous):

would you mind helping me with another sequence?

OpenStudy (freckles):

sure i can try

OpenStudy (anonymous):

(thanks!) 1, 1, 1/3, 1/4, 1/15...

OpenStudy (freckles):

i'm thinking

OpenStudy (freckles):

@Kainui @ganeshie8

OpenStudy (freckles):

I guess there are no hints?

OpenStudy (anonymous):

No

OpenStudy (kainui):

are there any limits on how many terms can be in the recurrence relation?

OpenStudy (anonymous):

It's an infinite sequence

OpenStudy (anonymous):

It just says to write a recursive rule for the sequence

OpenStudy (kainui):

Here's my recurrence relation: \[\frac{1}{3(a_0+a_1+3a_2+4a_3+1)}=a_4\]

ganeshie8 (ganeshie8):

\[a_n = \dfrac{a_{n-2}}{n}\\~\\a_1 = a_2=1\]

OpenStudy (freckles):

that's very groovy @ganeshie8

OpenStudy (freckles):

I don't think I would have seen that

OpenStudy (freckles):

like ever

OpenStudy (anonymous):

Ohhh, so that's how you do it. Thanks! Is there a way to calculate it or is it just guess and check?

ganeshie8 (ganeshie8):

I have cheated : https://oeis.org/A081405

ganeshie8 (ganeshie8):

There is no unique answer as we can cookup infinitely many sequences having the first few terms as common one boring, but systematic way is to find a polynomial fit of the given terms...

OpenStudy (anonymous):

Oh ok. Thank you very much for your help.